OpenStudy (anonymous):

simplify. -(-9) +2y + (-9) + 6

OpenStudy (compassionate):

= 9 + 2y - 9 + 6 = 2y + 6

OpenStudy (anonymous):

thats what i got.. just making sure.

OpenStudy (compassionate):

Okay, good luck!

OpenStudy (compassionate):

========================> Can I get bet answer, please :)

OpenStudy (anonymous):

could you help me with this one... express the answer to each question as the difference between two real numbers. compute the difference. answer the question and interpret the sign of the answer. the highest recorded weather temperature on earth is 58.0 celcious. the difference between the record and the lowest recorded weather temperature is 146.3 celceous. what is the record low?

OpenStudy (compassionate):

-88.3

OpenStudy (anonymous):

thank you! what about this one.. pythagoras, the greek mathematician, was born in 582 b.c. and died on his birthday in 497 b.c. how old was he when he died?

OpenStudy (compassionate):

85. These problems are actually very easy. Want me to explain them to you?

OpenStudy (anonymous):

yes please

OpenStudy (compassionate):

Okay, let me give you an example. What is your age?

OpenStudy (anonymous):

13

OpenStudy (compassionate):

So you were born in 1999.

OpenStudy (compassionate):

Correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

98

OpenStudy (compassionate):

Alright. Well, for 13, let's assume you were born in 1999. I can get your age by subtracting 2012 - 1999 = 13. This is known as the 'difference.' The difference between two numbers is determined by subtracting the smaller number from the larger number. For example the difference between 12 and 9 is: 12-9=3. If I said that I'm 23 years old, and you're 13, and I wanted the difference between our ages, all I would have to do is subtract 13 from 23 23 - 13 = 10, we're ten years apart. For example the difference between: 100 and 25 is 75 100 and -25 is 125 -100 and -115 is 15 -500 and 100 is 600

OpenStudy (anonymous):

oh okay thats helps.. so pretty much all im doing is subtracting

OpenStudy (compassionate):

Yea, basically.

OpenStudy (anonymous):

oh okay!